Evaldez enterprise invest 20,000 part at 12% and the rest 14%. How much should be invested in each rate so that the income from the 14% investment is half as much as the income from the 12% investment
Answer by mananth(16949)

12.00% per annum ------------- x
14.00% per annum ------------ (20,000-x)
total investment 20000
interest from 14.00% = 0.5times interest from 12.00%

14.00% *x = 0.5* 12.00% ( 20000 - x)
14.00% x = 1200 - 0.06 x
0.14 x = 1200 - 0.06 x
0.2 x = 1200
x= 6000 14000
12.00% per annum ------------- $6,000
14.00% per annum ------------ $14,000